OpenStudy (anonymous):

can somebody please help me with this because i keep getting 15.3 and that is wrong Ratio of area of circle C to its radius is 12 times pi . Which is the circumference of the circle

OpenStudy (anonymous):

\[A=\frac{\theta \times r^2}{2}\] So for a whole circle \[\theta=2 \pi\] You say \[A=12\pi r\] So \[12\pi r = \pi r^2\] \[r=12\] And circumference is given by \[2 \pi r = 2 \pi \times 12 = 24 \pi\]

OpenStudy (anonymous):

thnks